# Antinucleon-nucleon interactions in covariant chiral effective field theory

### v1

@@ Abstract
Motivated by the recent progress in developing high-precision relativistic chiral nucleon-nucleon interactions, we study the antinucleon-nucleon interaction at the leading order in the covariant chiral effective field theory. The phase shifts and inelasticities with $J\leq 1$ are obtained and compared to their non-relativistic counterparts. For most partial waves, the descriptions of phase shifts and inelasticities in the leading-order covariant chiral effective field theory are comparable to those in the next-to-leading order non-relativistic chiral effective field theory, confirming the relatively faster convergence observed in the nucleon-nucleon sector. In addition, we search for bound states/resonances near the $\bar{N}N$ threshold and find several structures that can be associated with those states recently observed by the BESIII Collaboration.
@@ I Introduction
There has been ongoing interest in antinucleon-nucleon ($\bar{N}N$) interactions over the last decade. One primary motivation is the observations of near-threshold $\bar{N}N$ enhancements in charmonium  decays Bai et al. (2003); Ablikim et al. (2005, 2012, 2016, 2022, 2024), $B$ meson decays Abe et al. (2002a, b), and $e^{+}e^{-}\rightarrow\bar{p}p$ reactions Aubert et al. (2005, 2006) . Those observations provided an opportunity to elucidate the existence of speculated $\bar{N}N$ molecules and stimulate studies of the $\bar{N}N$ interactions at low energies. Other motivations include the novel proposal of a super $J/\psi$ factory Yuan and Karliner (2021) and the construction of next-generation facilities, such as the Facility for Antiproton and Ion Research (FAIR) in Darmstadt Sturm et al. (2010) and the Super Tau-Charm Facility (STCF) in Huizhou Peng et al. (2020).
@@
In this work, we construct the antinucleon-nucleon interaction in covariant chiral effective field theory at leading order (LO). A relativistic three-dimensional reduction of the Bethe-Salpeter equation is used to obtain the scattering amplitude from the chiral potential. All 26 LECs parameterizing the short-range and annihilation potentials are fixed by fitting to the energy-dependent Nijmegen partial wave analysis (PWA) of the $\bar{p}p$ data Zhou and Timmermans (2012). A satisfactory description of the phase shifts and inelasticities of low angular momenta is achieved in analogy to the pertinent relativistic $NN$ interaction.
The paper is organized as follows. In Sect. II, we explain how to derive the leading-order chiral potentials. The scattering equation and the procedure to obtain the phase shifts are shown in Sect. III. In this formalism, the $\bar{N}N$ phase shifts for $J\leq 1$ partial waves are calculated, and possible near-threshold bound/resonant states are searched for in Sect. IV. Finally, we provide a summary in Sect. V.
II Chiral potentials at leading order
Figure 1: Feynman diagrams contributing to the $\bar{N}N$ interaction at leading order in the covariant power counting. The solid lines denote nucleons/antinucleons, and the dashed line represents the pion. The box denotes the vertex from $\mathcal{L}_{\bar{N}N}^{(0)}$, while the circle/dot shows vertex from $\mathcal{L}_{\pi\bar{N}}^{(1)}$/$\mathcal{L}_{\pi N}^{(1)}$.
The $\bar{N}N$ interaction contains scattering and annihilation potentials, which reads
$V_{\bar{N}N}=V^{\text{S}}+V^{\text{A}}.$ (1)
For the scattering process, the underlying covariant power counting of the $\bar{N}N$ interaction is the same as the $NN$ case, which is described in detail in Refs. Ren et al. (2018); Xiao et al. (2019); Lu et al. (2022), because the antinucleon field (spinor $v$) and the nucleon field (spinor $u$) are treated on an equal footing as spin-1/2 fields. The corresponding Feynman diagrams at LO are summarized in Fig. 1, and the relevant Lagrangians are,
@@ II Chiral potentials at leading order
The covariant scattering potentials at leading order $V^{\text{S}}_{\text{LO}}$ can be obtained by summing the contact (CT) and one-pion-exchange (OPE) terms shown in Fig. 1,
$V_{\text{LO}}^{\text{S}}=V^{\text{S}}_{\text{CT}}+V^{\text{S}}_{\text{OPE}},$ (10)
where the contact potential $V^{\text{S}}_{\text{CT}}$ is,
$V^{\text{S}}_{\text{CT}}\left(\bm{p},\bm{p}^{\prime}\right)=$ $C_{S}\left[\bar{v}\left(\bm{p},s_{1}\right)v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)u\left(-\bm{p},s_{2}\right)\right]+C_{A}\left[\bar{v}\left(\bm{p},s_{1}\right)\gamma_{5}v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)\gamma_{5}u\left(-\bm{p},s_{2}\right)\right]$ (11)
@@
and the one-pion-exchange potential $V^{\text{S}}_{\text{OPE}}$ is
$V^{\text{S}}_{\text{OPE}}\left(\bm{p},\bm{p}^{\prime}\right)=\frac{g_{A}^{2}}{4f_{\pi}^{2}}\frac{\left[\bar{v}\left(\bm{p},s_{1}\right)\bm{\tau}_{1}\gamma_{\mu}\gamma_{5}q^{\mu}v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)\bm{\tau}_{2}\gamma_{\nu}\gamma_{5}q^{\nu}u\left(-\bm{p},s_{2}\right)\right]}{\left(E_{p^{\prime}}-E_{p}\right)^{2}-\left(\bm{p}^{\prime}-\bm{p}\right)^{2}-m_{\pi}^{2}},$ (12)
@@
where $m_{N}$ refers to the nucleon mass, and we use the isospin-averaged value $m_{N}=939$ MeV, $\epsilon_{p}=E_{p}+m_{N}$, $\chi_{s}$ denotes the Pauli spinor matrix, and $\bm{\sigma}$ is the Pauli matrix. The Spinor $v\left(\bm{p},s\right)=\gamma_{0}Cu^{*}\left(\bm{p},s\right)$ with $C$ represents the charge transformation operator,
@@
where $\bm{e}\left(abcd\right)$ represents an ordering of quadrilinears and $\bm{K}^{\left(abcd\right)}$ is the transformation matrix, whose explicit forms are given in the Appendix A. This allows one to express all the arrangements as a linear combination of the chosen type, in our case, Eq. (11). Here the $C_{A}\left[\bar{v}\left(\bm{p},s_{1}\right)\gamma_{5}v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)\gamma_{5}u\left(-\bm{p},s_{2}\right)\right]$ term, which arises from the next-to-leading order potential according to Refs. Xiao et al. (2019); Lu et al. (2022), is ascended to leading order to ensure that one can make use of the generalized Fierz identities to get rid of redundant terms in the potential.
@@
$V^{\text{S}}_{1S0}=\xi\left[C_{1S0}\left(R_{p}^{2}+R_{p^{\prime}}^{2}\right)+\hat{C}_{1S0}\left(1+R_{p}^{2}R_{p^{\prime}}^{2}\right)\right],$
$V^{\text{S}}_{3P0}=\xi C_{3P0}R_{p}R_{p^{\prime}},$
$V^{\text{S}}_{1P1}=2\xi\left(C_{3S1}-3\hat{C}_{3S1}\right)R_{p}R_{p^{\prime}},$
$V^{\text{S}}_{3P1}=\frac{4}{3}\xi\left(C_{1S0}-\hat{C}_{1S0}\right)R_{p}R_{p^{\prime}},$
$V^{\text{S}}_{3S1}=\xi\left[C_{3S1}\left(R_{p}^{2}+R_{p^{\prime}}^{2}\right)+\hat{C}_{3S1}\left(9+R_{p}^{2}R_{p^{\prime}}^{2}\right)\right],$
$V^{\text{S}}_{3D1}=8\xi\hat{C}_{3S1}R_{p}^{2}R_{p^{\prime}}^{2},$
$V^{\text{S}}_{3S1-3D1}=2\sqrt{2}\xi\left(2C_{3S1}R_{p}^{2}+\hat{C}_{3S1}R_{p}^{2}R_{p^{\prime}}^{2}\right),$
$V^{\text{S}}_{3D1-3S1}=2\sqrt{2}\xi\left(2C_{3S1}R_{p^{\prime}}^{2}+\hat{C}_{3S1}R_{p}^{2}R_{p^{\prime}}^{2}\right),$ (16)
@@
Compared to the $NN$ interaction, a new feature of the $\bar{N}N$ interaction is the presence of the annihilation process, which leads to an intrinsic difficulty in describing a system that has hundreds of annihilation many-body channels at rest Carbonell et al. (2023). Here, we follow the approach of Ref. Kang et al. (2014) that manifestly fulfills unitarity and considers the annihilation potential of the following form,
$V^{\text{A}}=\sum_{X=2\pi,3\pi,...}V_{\bar{N}N\rightarrow X}G_{X}V_{X\rightarrow\bar{N}N},$ (18)
@@
$\text{Im}V^{\text{A}}=-\pi\sum_{X}V_{\bar{N}N\rightarrow X}V_{X\rightarrow\bar{N}N}.$ (20)
@@
Then, expanding $V_{\bar{N}N\rightarrow X}$ in powers of the nucleon three momentum up to next-to-leading order (NLO), one can obtain the annihilation potential,
$V_{1S0}^{\text{A}}=-i\left(C_{1S0}^{a}+\hat{C}_{1S0}^{a}\frac{p^{2}}{4m_{N}^{2}}\right)\left(C_{1S0}^{a}+\hat{C}_{1S0}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}}\right),$
$V_{3P0}^{\text{A}}=-i\left(C_{3P0}^{a}\right)^{2}\frac{pp^{\prime}}{4m_{N}^{2}},$
$V_{1P1}^{\text{A}}=-i\left(C_{1P1}^{a}\right)^{2}\frac{pp^{\prime}}{4m_{N}^{2}},$
$V_{3P1}^{\text{A}}=-i\left(C_{3P1}^{a}\right)^{2}\frac{pp^{\prime}}{4m_{N}^{2}},$
$V_{3S1}^{\text{A}}=-i\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{2}}{4m_{N}^{2}}\right)\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}}\right),$
$V_{3S1-3D1}^{\text{A}}=-i\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{2}}{4m_{N}^{2}}\right)C_{\epsilon_{1}}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}},$
$V_{3D1-3S1}^{\text{A}}=-iC_{\epsilon_{1}}^{a}\frac{p^{2}}{4m_{N}^{2}}\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}}\right),$
$V_{3D1}^{\text{A}}=-i\left(C_{\epsilon_{1}}^{a}\right)^{2}\frac{p^{2}p^{\prime 2}}{16m_{N}^{4}}.$ (21)
@@
The factors $\frac{1}{4m_{N}^{2}}$ and $\frac{1}{16m_{N}^{4}}$ are introduced to ensure that all annihilation constants are of the same dimension. There are several issues to address regarding the annihilation potential. 1) Eq. (II) only contains the $J=0,1$ partial waves to be consistent with the scattering potential given in Eq. (II). 2) Eq. (II) is organized in the conventional Weinberg power counting, while a more self-consistent annihilation potential should be constructed in the covariant power counting. The main difficulty in evaluating a covariant annihilation potential is the complexity of the explicit expressions for all open annihilation potentials $V_{\bar{N}N\rightarrow X}$. Based on the experience in the $NN$ interaction, the accuracy of the covariant potential is comparable to the non-relativistic potential at one order higher. Therefore, we expand $V_{\bar{N}N\rightarrow X}$ up to NLO in the conventional Weinberg power counting to evaluate the annihilation potential as an approximation of the exact covariant annihilation potential at LO. 3) The full annihilation potential contains a real part from the principal value in Eq. (18), whose structure is accounted for by the LECs in the conventional non-relativistic scattering potential at the corresponding order. By contrast, the contribution of the real part of Eq. (18) can only be absorbed in the covariant LECs partly in our case because the number of independent LECs in the covariant power counting at LO is less than that in the non-relativistic power counting at NLO. However, this real part does not break unitarity. In addition, the contribution to the $\bar{N}N$ interaction from the additional structures is suppressed because it is of order $\mathcal{O}\left(p^{2}\right)$ (here $p$ refers to the small quantity in the conventional Weinberg power counting). Therefore, we use the pure imaginary potential in Eq. (II) to account for the annihilation process in practice. It should be mentioned that the problem above can be solved by constructing a self-consistent annihilation potential in the covariant power counting. We will explore how to implement this idea in the future.

### v2

# Reexamination of antinucleon-nucleon interactions in covariant chiral effective field theory

@@ Abstract
Motivated by the recent progress in developing high-precision relativistic chiral nucleon-nucleon interactions, we study the antinucleon-nucleon interaction in a hybrid approach where the real part of the potential is constructed in the leading-order covariant chiral effective field theory, and the imaginary part is described following the procedure adopted in the heavy baryon chiral effective field theory. The phase shifts and inelasticities with $J\leq 1$ are obtained and compared to those calculated in the next-to-leading order heavy baryon chiral effective field theory. For most partial waves, the descriptions of phase shifts and inelasticities in the hybrid approach are comparable to those in the next-to-leading order heavy baryon chiral effective field theory, confirming the relatively faster convergence of the relativistic approach observed in the nucleon-nucleon sector. In addition, we search for bound states/resonances near the $\bar{N}N$ threshold and find several structures that can be associated with those states recently observed by the BESIII Collaboration.
@@ I Introduction
There has been ongoing interest in antinucleon-nucleon ($\bar{N}N$) interactions over the last decade. One primary motivation is the observations of near-threshold $\bar{N}N$ enhancements in charmonium  decays Bai et al. (2003); Ablikim et al. (2005, 2012, 2016, 2022, 2024), $B$ meson decays Abe et al. (2002a, b), and $e^{+}e^{-}\rightarrow\bar{p}p$ reactions Aubert et al. (2005, 2006) . Those observations provided an opportunity to elucidate the existence of speculated $\bar{N}N$ molecules and stimulated studies of the $\bar{N}N$ interactions at low energies. Other motivations include the novel proposal of a super $J/\psi$ factory Yuan and Karliner (2021) and the construction of next-generation facilities, such as the Facility for Antiproton and Ion Research (FAIR) in Darmstadt Sturm et al. (2010) and the Super Tau-Charm Facility (STCF) in Huizhou Peng et al. (2020).
@@
In this work, we construct the antinucleon-nucleon interaction in a hybrid approach in which the elastic process is determined in the leading order (LO) covariant ChEFT, and the annihilation process is described in the next-to-leading order (NLO) heavy baryon (HB) ChEFT. A relativistic three-dimensional reduction of the Bethe-Salpeter equation is used to obtain the scattering amplitude from the hybrid potential. All 26 LECs parameterizing the short-range and annihilation potentials are fixed by fitting to the energy-dependent Nijmegen partial wave analysis (PWA) of the $\bar{p}p$ data Zhou and Timmermans (2012). A satisfactory description of the phase shifts and inelasticities of low angular momenta is achieved in analogy to the pertinent relativistic $NN$ interaction.
The paper is organized as follows. In Sect. II, we explain how to derive the hybrid potentials. The scattering equation and the procedure to obtain the phase shifts are shown in Sect. III. In Sect. IV, the $\bar{N}N$ phase shifts for $J\leq 1$ partial waves are calculated, and possible near-threshold bound/resonant states are searched for. Finally, we provide a summary in Sect. V.
II The hybrid antinucleon-nucleon potentials
Figure 1: Feynman diagrams contributing to the elastic part of the $\bar{N}N$ interaction at leading order in the covariant power counting. The solid lines denote nucleons/antinucleons, and the dashed line represents the pion. The box denotes the vertex from $\mathcal{L}_{\bar{N}N}^{(0)}$, while the circle/dot shows vertex from $\mathcal{L}_{\pi\bar{N}}^{(1)}$/$\mathcal{L}_{\pi N}^{(1)}$.
The $\bar{N}N$ interaction contains real and imaginary parts, which reads
$V_{\bar{N}N}=V^{\text{R}}+V^{\text{I}}.$ (1)
For the real parts of the interaction, the underlying covariant power counting is the same as the $NN$ case, which is described in detail in Refs. Ren et al. (2018); Xiao et al. (2019); Lu et al. (2022), where the antinucleon field (spinor $v$) and the nucleon field (spinor $u$) are treated on an equal footing as spin-1/2 fields. The corresponding LO Feynman diagrams are summarized in Fig. 1, and the relevant Lagrangians are,
@@ II The hybrid antinucleon-nucleon potentials
The LO covariant potentials $V^{\text{R}}_{\text{LO}}$ can be obtained by summing the contact (CT) and one-pion-exchange (OPE) terms shown in Fig. 1,
$V_{\text{LO}}^{\text{R}}=V^{\text{R}}_{\text{CT}}+V^{\text{R}}_{\text{OPE}},$ (10)
where the contact potential $V^{\text{R}}_{\text{CT}}$ is,
$V^{\text{R}}_{\text{CT}}\left(\bm{p},\bm{p}^{\prime}\right)=$ $C_{S}\left[\bar{v}\left(\bm{p},s_{1}\right)v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)u\left(-\bm{p},s_{2}\right)\right]+C_{A}\left[\bar{v}\left(\bm{p},s_{1}\right)\gamma_{5}v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)\gamma_{5}u\left(-\bm{p},s_{2}\right)\right]$ (11)
@@
and the one-pion-exchange potential $V^{\text{R}}_{\text{OPE}}$ is
$V^{\text{R}}_{\text{OPE}}\left(\bm{p},\bm{p}^{\prime}\right)=\frac{g_{A}^{2}}{4f_{\pi}^{2}}\frac{\left[\bar{v}\left(\bm{p},s_{1}\right)\bm{\tau}_{1}\gamma_{\mu}\gamma_{5}q^{\mu}v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)\bm{\tau}_{2}\gamma_{\nu}\gamma_{5}q^{\nu}u\left(-\bm{p},s_{2}\right)\right]}{\left(E_{p^{\prime}}-E_{p}\right)^{2}-\left(\bm{p}^{\prime}-\bm{p}\right)^{2}-m_{\pi}^{2}},$ (12)
@@
where $m_{N}$ refers to the nucleon mass, and we use the isospin-averaged value $m_{N}=939$ MeV, $\epsilon_{p}=E_{p}+m_{N}$, $\chi_{s}$ denotes the Pauli spinor matrix, and $\bm{\sigma}$ is the Pauli matrix. The Spinor $v\left(\bm{p},s\right)=\gamma_{0}Cu^{*}\left(\bm{p},s\right)$ with $C$ representing the charge transformation operator,
@@
where $\bm{e}\left(abcd\right)$ represents an ordering of quadrilinears and $\bm{K}^{\left(abcd\right)}$ is the transformation matrix, whose explicit forms are given in Appendix A. This allows one to express all the arrangements as a linear combination of the chosen type, in our case, Eq. (11). Here the $C_{A}\left[\bar{v}\left(\bm{p},s_{1}\right)\gamma_{5}v\left(\bm{p}^{\prime},s_{1}^{\prime}\right)\right]\left[\bar{u}\left(-\bm{p}^{\prime},s_{2}^{\prime}\right)\gamma_{5}u\left(-\bm{p},s_{2}\right)\right]$ term, which arises from the next-to-leading order potential according to Refs. Xiao et al. (2019); Lu et al. (2022), is ascended to leading order to ensure that one can make use of the generalized Fierz identities to get rid of redundant terms in the potential.
@@
$V^{\text{R}}_{1S0}=\xi\left[C_{1S0}\left(R_{p}^{2}+R_{p^{\prime}}^{2}\right)+\hat{C}_{1S0}\left(1+R_{p}^{2}R_{p^{\prime}}^{2}\right)\right],$
$V^{\text{R}}_{3P0}=\xi C_{3P0}R_{p}R_{p^{\prime}},$
$V^{\text{R}}_{1P1}=2\xi\left(C_{3S1}-3\hat{C}_{3S1}\right)R_{p}R_{p^{\prime}},$
$V^{\text{R}}_{3P1}=\frac{4}{3}\xi\left(C_{1S0}-\hat{C}_{1S0}\right)R_{p}R_{p^{\prime}},$
$V^{\text{R}}_{3S1}=\xi\left[C_{3S1}\left(R_{p}^{2}+R_{p^{\prime}}^{2}\right)+\hat{C}_{3S1}\left(9+R_{p}^{2}R_{p^{\prime}}^{2}\right)\right],$
$V^{\text{R}}_{3D1}=8\xi\hat{C}_{3S1}R_{p}^{2}R_{p^{\prime}}^{2},$
$V^{\text{R}}_{3S1-3D1}=2\sqrt{2}\xi\left(2C_{3S1}R_{p}^{2}+\hat{C}_{3S1}R_{p}^{2}R_{p^{\prime}}^{2}\right),$
$V^{\text{R}}_{3D1-3S1}=2\sqrt{2}\xi\left(2C_{3S1}R_{p^{\prime}}^{2}+\hat{C}_{3S1}R_{p}^{2}R_{p^{\prime}}^{2}\right),$ (16)
@@
Compared to the $NN$ interaction, a new feature of the $\bar{N}N$ interaction is the presence of the annihilation process, which leads to an intrinsic difficulty in describing a system that has hundreds of annihilation many-body channels at rest Carbonell et al. (2023). Here, we follow the approach of Ref. Kang et al. (2014) that manifestly fulfills unitarity and considers the contributions to the potential from the annihilation process of the following form,
$V=\sum_{X=2\pi,3\pi,...}V_{\bar{N}N\rightarrow X}G_{X}V_{X\rightarrow\bar{N}N},$ (18)
@@
$\text{Im}V=-\pi\sum_{X}V_{\bar{N}N\rightarrow X}V_{X\rightarrow\bar{N}N}.$ (20)
@@
Expanding $V_{\bar{N}N\rightarrow X}$ in powers of the nucleon three momentum up to NLO , one can obtain the imaginary parts of the $\bar{N}N$ interaction,
$V_{1S0}^{\text{I}}=-i\left(C_{1S0}^{a}+\hat{C}_{1S0}^{a}\frac{p^{2}}{4m_{N}^{2}}\right)\left(C_{1S0}^{a}+\hat{C}_{1S0}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}}\right),$
$V_{3P0}^{\text{I}}=-i\left(C_{3P0}^{a}\right)^{2}\frac{pp^{\prime}}{4m_{N}^{2}},$
$V_{1P1}^{\text{I}}=-i\left(C_{1P1}^{a}\right)^{2}\frac{pp^{\prime}}{4m_{N}^{2}},$
$V_{3P1}^{\text{I}}=-i\left(C_{3P1}^{a}\right)^{2}\frac{pp^{\prime}}{4m_{N}^{2}},$
$V_{3S1}^{\text{I}}=-i\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{2}}{4m_{N}^{2}}\right)\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}}\right),$
$V_{3S1-3D1}^{\text{I}}=-i\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{2}}{4m_{N}^{2}}\right)C_{\epsilon_{1}}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}},$
$V_{3D1-3S1}^{\text{I}}=-iC_{\epsilon_{1}}^{a}\frac{p^{2}}{4m_{N}^{2}}\left(C_{3S1}^{a}+\hat{C}_{3S1}^{a}\frac{p^{\prime 2}}{4m_{N}^{2}}\right),$
$V_{3D1}^{\text{I}}=-i\left(C_{\epsilon_{1}}^{a}\right)^{2}\frac{p^{2}p^{\prime 2}}{16m_{N}^{4}}.$ (21)
The factors $\frac{1}{4m_{N}^{2}}$ and $\frac{1}{16m_{N}^{4}}$ are introduced to ensure that all annihilation constants are of the same dimension.
@@
There are several issues to address regarding the contributions to the potential from the annihilation process. 1) Eq. (II) only contains the $J=0,1$ partial waves to be consistent with the real parts of the potential given in Eq. (II). 2) Eq. (II) is organized in the conventional Weinberg power counting, while a more self-consistent potential should be constructed in the covariant power counting. The main difficulty in evaluating a covariant potential relevant to the annihilation process is the complexity of the explicit expressions for all open annihilation channels’ potentials $V_{\bar{N}N\rightarrow X}$. Based on the experience in the $NN$ interaction, the accuracy of the covariant potential is comparable to the HB potential at one order higher. Therefore, we expand $V_{\bar{N}N\rightarrow X}$ up to NLO in the conventional Weinberg power counting to evaluate the imaginary parts of the potential as an approximation of the exact covariant potential at LO. 3) The complete contributions to the potential from the annihilation process contain a real part from the principal value in Eq. (18), whose structure is accounted for by the real parts of the LECs in the conventional chiral potential at the corresponding order. By contrast, the contribution of the real part of Eq. (18) can only be absorbed in the covariant LECs partly in our case because the number of independent LECs in the covariant power counting at LO is less than that in the conventional Weinberg power counting at NLO. However, this real part does not break unitarity. In addition, the contribution to the $\bar{N}N$ interaction from the additional structures is suppressed because it is of order $\mathcal{O}\left(p^{2}\right)$ (here $p$ refers to the small quantity in the conventional Weinberg power counting). Therefore, we use the pure imaginary potential in Eq. (II) to account for the annihilation process in practice. It should be mentioned that the problem above can be solved by constructing a self-consistent potential responsible for the annihilation process in the covariant power counting. We will explore how to implement this idea in the future.